OpenStudy (puppylover52):

A porch has the shape of a semicircle with a diameter of 7 ft. What is the approximate area of the porch, rounded to the nearest square foot? Use 3.14 to approximate .

OpenStudy (puppylover52):

@fairytail_lover @YoungStudier @haleyright @jamye

OpenStudy (puppylover52):

I dont need the answer i need help

rebeccaxhawaii (rebeccaxhawaii):

okay well first get your radius which is half the diameter so 7/2=3.5

OpenStudy (puppylover52):

where did you get 2?

rebeccaxhawaii (rebeccaxhawaii):

the radius is half the diameter therefore 7/2 we are splitting it in half

OpenStudy (puppylover52):

oh ok

rebeccaxhawaii (rebeccaxhawaii):



rebeccaxhawaii (rebeccaxhawaii):

the formula to find area for a circle is pi*r^2 so 3.14*3.5^2

OpenStudy (puppylover52):

i think its 5.495???

rebeccaxhawaii (rebeccaxhawaii):

\[3.5^2=3.5*3.5\]

rebeccaxhawaii (rebeccaxhawaii):

what is 3.5*3.5=?

OpenStudy (puppylover52):

10.99

rebeccaxhawaii (rebeccaxhawaii):

please type it into your calculator. 3.5*3.5

OpenStudy (puppylover52):

I did

OpenStudy (puppylover52):

nvm

rebeccaxhawaii (rebeccaxhawaii):

try again

OpenStudy (puppylover52):

12.25 sorry i read it wrong

rebeccaxhawaii (rebeccaxhawaii):

no no its okay. now multiply that by pi which is 3.14 so 12.25*3.14=?

OpenStudy (puppylover52):

38.465

rebeccaxhawaii (rebeccaxhawaii):

great so you round to the nearest square foot which is 38